OpenStudy (anonymous):

The area of this cylinder's base is 150 square inches. The height of the cylinder is 6 inches. What is the volume of the cone that has the same base area and height? Show the steps of your solution and explain your work

OpenStudy (anonymous):

\[A_{cylinder base}=\pi r^2=150\] \[A{cone}=\frac{1}{3}\pi r h\]

OpenStudy (anonymous):

find r from equation (1)

OpenStudy (anonymous):

whats the volume of the cone thou

OpenStudy (anonymous):

that has same base area and hieght

OpenStudy (anonymous):

\[A_{cone}=\frac{1}{3}\pi r^2h\] \[\pi r^2=150 \implies \frac{ 1 }{ 3 }\color{red}{\pi r^2}h=\frac{1}{3}(\color{red}{150})h\] h=6

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

i got 2700 as volume

OpenStudy (anonymous):

is it not 1/3*150*6=300